Simplify: (-3a^2b^-2)^2 / (5a^3)^2
Thanks guys for your help, this is a huge help

Answer by vasu2qute(17)

Solution :
(-3a^2b^-2)^2 / (5a^3)^2=
(-3^2 x a^4 x b^-4)/(5^2 x a^6) =
(-3 x -3 x a^4 x b^-4))/(5 x 5 x a^6) =
9a^4b^-4/25a^6 = (9a^-2b^-4)/25